# Guaranteed Generation from Large Language Models

### v1

@@ 4.2 Conditional generation with a positive ending constraint
First, we prepared a positive sentiment scorer $b_{score}(y)$ [9] and defined our binary, threshold-based constraint as $b_{\tau}(y)=1$ if and only if $b_{score}(y)>\tau$. We then selected our set of openings $X$ by collecting negative story openings from the ROCStories test set [29], using only the first two sentences of stories for $X$ and imposing that $b_{score}(X)<0.05$. To incorporate the domain knowledge of ROCStories, we use GPT-2 [35] fine-tuned on the ROCStories training set as our base LLM $a$. In this task, given a negative opening $X$, a completion $y=[Y;Z]$ is accepted if its final sentence $Z$ satisfies $b_{score}(Z)>0.98$. The acceptance rate for the base LLM $a$ is $AR_{a}$ = 0.005; in other words, rejection sampling from $a$ approximately incurs a 200$\times$ reduction in the acceptance rate compared to the ideal sampler $\tilde{g}$.
Empirical results
As shown in Figure 6, constraint-aware prompting improves the acceptance rate without any training, but results in a distribution that greatly diverges from $g$. Due to GPT-2 being a smaller model than Gemma-2B, the AR improvement from prompting is relatively less pronounced than in the lexical constraint experiment. In contrast, SFT and DPG, which optimize $\text{KL}(g||a^{\prime})$, demonstrate an increase in AR while simultaneously showing less divergence from the target distribution. DPG, in particular, exhibits a distribution closer to the gold model $g$. In summary, the GUARD algorithm yields an AR improvement of around 60$\times$ compared to $a$ while preserving a high proximity to $g$.
Analysis
@@
The results reported in Table 2 indicate that when conditioning on a specific prefix, the diversity is relatively lower compared to unconditional generation (as seen in Table 1). Nonetheless, samples from DPG-based $g^{\prime}$ and $g$ exhibit similar diversity levels. Instead, samples obtained through CAP resulted in higher similarity, i.e., lower diversity. Figure 7 demonstrates that a positional bias is also observed in the case of a sentiment constraint. Compared to $g$ and DPG-based $g^{\prime}$, when prompting to include a positive ending, the story continuation shows a more abrupt early shift towards positivity. This further confirms the existence of a positional bias in the constraint-aware prompting behavior — similarly to the lexical constraint scenario — and illustrates how this bias leads to distributional distortion. We provide additional details and results for this setting in App. E , including examples of generated samples.
@@ 5 Related work
LLM alignment techniques are widely regarded as the de facto method for integrating external desirable properties (e.g., helpfulness or harmlessness) into base language models. These techniques require access to a preference dataset, which consists of prompts, each associated with several completions, one of which is labeled as preferable based on criteria like human judgment. Popular approaches in this family include RLHF [32], which trains a reward model from the preference dataset and fine-tunes the LLM through PPO [40], and DPO [36], which combines these steps to directly fine-tune the model on the preference dataset. While these methods enable models to generate outputs that are better aligned with the implicit properties and values expressed in the preference dataset, such as harmlessness, there remains a gap between alignment and the goal of ensuring that all outputs are harmless. Alignment does not provide a straightforward means to enforce explicit constraints, as required in our guaranteed generation task. Consequently, despite their effectiveness, these alignment techniques, when used in a stand-alone manner (i.e., without additional inference-time processing), do not guarantee the satisfaction of a specific property or constraint in all generated outputs.
@@
More closely related to our goal of guaranteed generation is the line of work on controlled text generation, which explicitly defines conditions or constraints (e.g., in the form of predicates) that outputs should satisfy. This topic gathers a large body of literature, and we thus let the reader refer to Zhang et al. [46] for a more comprehensive survey. Existing approaches include prompting to integrate the constraint [45, 42], techniques related to reinforcement learning [13, 16, 24], or relying on inference-time methods to impose the conditions without updating the original language model [6, 17, 41, 22, 27, 14, 30, 4]. Although these methods allow for the explicit definition of constraints, as required in guaranteed generation, they share the same limitation as alignment techniques: they cannot enforce strict constraint satisfaction.
@@ 6 Conclusion and discussion
Appendix A Additional background and theoretical details
A.1 Characterization of $g$
@@
Let’s provide a self-contained proof of that fact. We have:
@@
A.2 On the limitations of autoregressive models in representing $g$
@@
First, we come back to the situation described in the main text where the constraint $b$ prohibits the token “bad” to appear in $y$. We saw that fine-tuning could not completely guarantee this constraint, because it would never put zero mass on the token “bad”. Furthermore, fine-tuning would be pretty inefficient, even as an approximator, for this “avoidance” situation: a huge dataset would probably be necessary for significantly reducing the probability of “bad”, because the fine-tuning process would only see examples not containing “bad”, with no explicit negative examples. As mentioned in the text, another approach, not involving fine-tuning, would be, at each timestep $t$ during the generation of $y$, to remove the token “bad” from the softmax vector $p(\cdot|y_{<t})$, and renormalize this vector to sum to 1. This would result in an autoregressive model $a^{\prime}$ that always respects the constraint, but whose distribution would differ from $g$. Intuitively, this would be because the distribution for $g$, which corresponds to natural sequences from $a$ not containing “bad”, would anticipate early on, for the initial tokens of $y$, the fact that “bad” cannot appear later, while $a^{\prime}$ would ignore any such influence. [^10] The intuition may be clearer with a more situation-specific word such as “malaria”: in order to eliminate sequences containing this word from $a$, $g$ has to depreciate the probability of prefixes that naturally lead to that word — such as “This proliferation of mosquitoes raises the risk of” — while an $a^{\prime}$ intervening on the local generation of “malaria” would not perform such anticipation.
@@
Conversely, rather than trying to avoid a certain token such as “bad”, we could try to enforce the presence of a token such as “good” somewhere in the sequence $y$. In other words, we would have $b(y)=1$ iff $y$ contains the token “good”. Here, we could again sample a large dataset from $a$, keep only the sequences containing “good”, and fine-tune an ARM $a^{\prime}$ on the resulting dataset. However, while $a^{\prime}$ would be better than $a$ at producing sequences containing “good”, it could still produce other sequences, and therefore would not strictly guarantee the constraint (although this fine-tuning would probably be more effective as an approximator than in the previous example of avoiding “bad”, because at least it could rely on instances explicitly containing the token). Again, we could try another autoregressive approach: we could, at each step of the generation of $y$, check whether we have already produced the token “good”, and if not, continue the generation of $y$ (i.e., prevent the generation of an $\langle eos\rangle$ token). This would result in an ARM $a^{\prime}$, which would guarantee strict satisfaction of the constraint, but the distribution of the sequences produced by $a^{\prime}$ would be very different from $g$. This is because, similarly to the previous case, while $g$ would anticipate in its distribution over early tokens the necessity to produce “good” at a later stage, $a^{\prime}$ would totally ignore such influence of “good” over the early tokens. Experiments could be conducted to confirm this fact, with the intuitive expectation that $a^{\prime}$ would tend to produce longer sequences than $g$, because there would be little incentive to produce “good” early on during the generation. [^11] Again, the intuition might be clearer if the constraint was about producing a rarer word than “good”, associated with a specific situation, for instance “trophy”. Sequences from $a$ containing this word, which would be selected by $g$, would tend to start differently than generic sequences from $a$, but $a^{\prime}$ would not show this kind of influence.
@@
For these two examples, the situation with standard LLMs such as $a$, built over neural networks, contrasts strongly with what would be the case if $a$ were an $n$-gram language model. In that case, $a$ would be equivalent to a probabilistic weighted finite-state automaton [28], a classical form of autoregressive model, and such constraints as the presence or absence of a specific token could be realized through intersection between such automata, resulting in another probabilistic automaton, with the same distribution as $g$. However, such intersections, which rely on dynamic programming over the finite set of states of these models, are unavailable in the domain of neural LMs. Thus, to summarize, if $a$ were an $n$-gram language model, where $b$ was a constraint similar to those considered in the two examples above, the problem of modeling $g$ through an ARM would be solvable. Another instance where this problem is solvable — staying this time in the familiar domain of neural language models — is when $a$ is a standard LLM, and $b$ is the constraint which imposes the presence of a certain prefix $[y_{1};\ldots;y_{t}]$ at the beginning of the generated sequence $y$. It is easy to see that the generative process that samples $y_{t+1}\sim a(\cdot|[y_{1};\ldots;y_{t}])$, $y_{t+2}\sim a(\cdot|[y_{1};\ldots;y_{t+1}])$, and so on, in the standard way, corresponds to an autoregressive model $a^{\prime}$ that has the same distribution as $g$. However, again, this is a very special case, and as soon as the constraint does not refer to such an initial section of $y$, the problem becomes much more challenging.
@@
A.3 Details about Theorem 1
@@
There exists a polynomial-time computable (PTC) autoregressive model $a$ and a PTC binary predicate $b$ such that no PTC autoregressive model $a^{\prime}$ has the same support [^12] The support of a distribution $p$ over ${\mathcal{Y}}$ is the set of $y$’s such that $p(y)>0$. as the gold model $g$, and in particular, has the same distribution as $g$.
@@ A.3.1 Proof sketch
[…]
@@
A.4 Characterization of $g^{\prime}$
Let $P^{\prime}(y)\doteq a^{\prime}(y)b(y)$ and $p^{\prime}(y)$ be the distribution such that $p^{\prime}(y)\propto P^{\prime}(y)$. Then $p^{\prime}(y)=\frac{1}{Z^{\prime}}P^{\prime}(y)$ with $Z^{\prime}\doteq\sum_{y\in{\mathcal{Y}}}P^{\prime}(y)$. We have $P^{\prime}(y)\leq a^{\prime}(y)\cdot 1$, and $r(y)\doteq\frac{P^{\prime}(y)}{a^{\prime}(y)\cdot 1}=b(y)$. According to the fundamental result about rejection sampling, see [23, Section 2.2], $p^{\prime}$ is the distribution obtained by sampling $y\sim a^{\prime}$ and accepting $y$ with probability $r(y)$, which in our case is either 00 or $1$. This is exactly what Algorithm 1 does, and therefore $g^{\prime}$ has the same distribution as $p^{\prime}$, in other words:

### v2

@@ 4.2 Conditional generation with a positive ending constraint
First, we prepared a positive sentiment scorer $b_{score}(y)$ [9] and defined our binary, threshold-based constraint as $b_{\tau}(y)=1$ if and only if $b_{score}(y)>\tau$. We then selected our set of openings $X$ by collecting negative story openings from the ROCStories test set [33], using only the first two sentences of stories for $X$ and imposing that $b_{score}(X)<0.05$. To incorporate the domain knowledge of ROCStories, we use GPT-2 [39] fine-tuned on the ROCStories training set as our base LLM $a$. In this task, given a negative opening $X$, a completion $y=[Y;Z]$ is accepted if its final sentence $Z$ satisfies $b_{score}(Z)>0.98$. The acceptance rate for the base LLM $a$ is $AR_{a}$ = 0.005 (i.e., on average, one sample is accepted every 200 samples drawn). In other words, rejection sampling from $a$ approximately incurs a 200$\times$ reduction in the acceptance rate compared to the ideal sampler $\tilde{g}$.
@@
Empirical results
As shown in Figure 7, constraint-aware prompting improves the acceptance rate without any training, but results in a distribution that greatly diverges from $g$. Due to GPT-2 being a smaller model than Gemma-2B, the AR improvement from prompting is relatively less pronounced than in the lexical constraint experiment. In contrast, SFT and DPG, which optimize $\text{KL}(g||a^{\prime})$, demonstrate an increase in AR while simultaneously showing less divergence from the target distribution. DPG, in particular, exhibits a distribution closer to the gold model $g$. In summary, the GUARD algorithm yields an AR improvement from 0.005 to 0.306 (almost 60$\times$) compared to $a$ while preserving a high proximity to $g$.
Analysis
@@
The results reported in Table 2 indicate that when conditioning on a specific prefix, the diversity is relatively lower compared to unconditional generation (as seen in Table 1). Nonetheless, samples from DPG-based $g^{\prime}$ and $g$ exhibit similar diversity levels. Instead, samples obtained through CAP resulted in higher similarity, i.e., lower diversity. Figure 6 demonstrates that a positional bias is also observed in the case of a sentiment constraint. Compared to $g$ and DPG-based $g^{\prime}$, when prompting to include a positive ending, the story continuation shows a more abrupt early shift towards positivity. This further confirms the existence of a positional bias in the constraint-aware prompting behavior — similarly to the lexical constraint scenario — and illustrates how this bias leads to distributional distortion. We provide additional details and results for this setting in App. F , including examples of generated samples.
@@ 5 Related Work
LLM alignment techniques are widely regarded as the de facto method for integrating external desirable properties (e.g., helpfulness or harmlessness) into base language models [15]. These techniques require access to a preference dataset, which consists of prompts, each associated with several completions, one of which is labeled as preferable based on criteria like human judgment. Popular approaches in this family include RLHF [36], which trains a reward model from the preference dataset and fine-tunes the LLM through PPO [44], and DPO [40], which combines these steps to directly fine-tune the model on the preference dataset. While these methods enable models to generate outputs that are better aligned with the implicit properties and values expressed in the preference dataset, such as harmlessness, there remains a gap between alignment and the goal of ensuring that all outputs are harmless. Alignment does not provide a straightforward means to enforce explicit constraints, as required in our guaranteed generation task. Consequently, despite their effectiveness, these alignment techniques, when used in a stand-alone manner (i.e., without additional inference-time processing), do not guarantee the satisfaction of a specific property or constraint in all generated outputs.
@@
More closely related to our goal of guaranteed generation is the line of work on controlled text generation, which explicitly defines conditions or constraints (e.g., in the form of predicates) that outputs should satisfy. This topic gathers a large body of literature, and we thus let the reader refer to Zhang et al. [53] for a more comprehensive survey. Existing approaches include prompting to integrate the constraint [50, 47], training techniques related to reinforcement learning [13, 17, 8, 26] or learning local discriminators to encourage constraint satisfaction [29], or else relying on inference-time methods to impose the conditions without updating the original language model [6, 18, 46, 27, 24, 7, 31, 14, 34, 4].
The previously cited papers do not focus on strictly enforcing constraint satisfaction. However, two recent works [51, 52] address this issue. Similarly to us, they are able to produce outputs that strictly satisfy the constraint, while maintaining proximity to the original model. They do so by approximating the LLM with a Hidden Markov Model (HMM), and by exploiting the dynamic programming properties of HMMs. This approach enables the computation of a weighted intersection of the HMM with a constraint over lexical items, resulting in a new HMM that always respects the constraint. Two key differences between their approach and ours are that (i) we handle arbitrary constraints — not necessarily of a lexical nature — which would be difficult to do through HMMs or similar finite-state mechanisms, and (ii) our focus is on producing an efficient sampler that is distributionally close to the gold target distribution $g$ that intrinsically represents the constraint $b$. In contrast, their focus is on producing a decoder that performs well on downstream tasks that have a looser relation to the constraint. We provide more details on these and some other related works in App. A.
@@ 6 Conclusion and Discussion
Acknowledgments
We would like to thank Laurent Besacier and Germán Kruszewski for their valuable feedback on this work.
MK acknowledges that this work was partially supported by an Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].
Appendix A Complements on Related Work
Relation of our work to Zhang et al. [51, 52]
Similar to us, the approach from Zhang et al. [51] is able to produce outputs that always satisfy the constraint, while maintaining proximity to the original model. It does so by approximating the LLM with an HMM, which, in contrast to the LLM, supports a dynamic programming approach to complying with lexical constraints over the output, and therefore the ability to weigh the long-term consequences of local next-token choices relative to these constraints. When the HMM approximation to the LLM is good enough (which depends in particular on the number of states of the HMM), the same HMM can be used at inference time with different lexical constraints without need of retraining.
While such an approach is relevant for guaranteed generation, we note several fundamental differences with what we do:
  1. 1.
    We handle arbitrary logical constraints (that is, binary predicates), not necessarily of a lexical nature, which would be difficult to handle through HMMs or similar finite-state mechanisms.
  2. 2.
    We give a central status to the “gold” constrained distribution $g$, which is the distribution that minimally deviates from the original distribution while still fully satisfying the constraint, and we evaluate the quality of our sampler in terms of distributional distance to this distribution $g$. Zhang et al. [51] do not focus on the evaluation of a sampler relative to a reference distribution, but rather on the evaluation of a decoder in terms of downstream tasks which have a looser relation with the constraint.
  3. 3.
    We study the trade-off between the quality of the sampler and its efficiency in the context of a simple rejection sampling mechanism based on an autoregressive approximation $a\textquoteright$ ’ of $g$ and show that quality and efficiency are both directly controlled by the divergence $\text{KL}(g||a\textquoteright)$ ’ ). This in turns motivates our interest in an approximation technique focused on minimizing this divergence, such as DPG.
The work from Zhang et al. [51] is extended in a follow-up [52] which considers lexical constraints defined through deterministic finite automata. As a side observation, in App. B.2 (Solvable instances of the problem), we have briefly mentioned that in case the base ARM and the lexical constraint are based on weighted finite state automata, the intersection of these automata could be constructed and sampled from. We note that such weighted automata are naturally very much related to HMMs, and might be interesting to study in their own right.
NADO sampler [29]
The NADO distribution introduced in Meng et al. [29] is trained with an objective similar to ours — namely, it tries to approximate a distribution comparable to our gold filtered distribution $g$. As a training method to obtain an autoregressive distribution similar to our $a\textquoteright$ ’, it has some analogies with (i) SFT, by training on samples from the filtered distribution, and (ii) DPG, by also performing a kind of distributional matching, but with more emphasis on local (i.e., token-level) decisions. Contrary to DPG, NADO does not directly have the objective of minimizing the divergence $\text{KL}(g||a\textquoteright)$ ’ ), which is the determining factor for the success of GUARD as shown in Theorem 2. Using NADO for $a\textquoteright$ ’ would nonetheless be interesting as a follow-up work, to see whether the value of its divergence is competitive with the cases we have explored, leading to improved performance for GUARD.
Appendix B Additional Background and Theoretical Details
B.1 Characterization of $g$
@@
Controlling the distortion of $a$’s capabilities through the use of $g$
The fact that $g(y)=\frac{1}{Z}a(y)\,b(y)$ implies that for any two sequences $y,y^{\prime}$ with $b(y)=b(y^{\prime})=1$ and $a(y)>0$, we have $g(y)>0$ and:
$\frac{g(y^{\prime})}{g(y)}=\frac{a(y^{\prime})}{a(y)}.$
In words, this means that if $y^{\prime}$ is $\alpha$ times more likely than $y$ relative to the original model $a$, and if both $y$ and $y^{\prime}$ respect the constraint, than $y^{\prime}$ is still exactly $\alpha$ times more likely than $y$ relative to $g$.
For example, suppose that, relative to $a$, the sentence “Potatoes were first cultivated in South America" is 100 times more likely than the sentence “Potatoes were first cultivated in Africa", and that both sentences satisfy the constraint, then the first sentence will still be 100 times more likely than the second relative to $g$.
So $g$ does not distort the knowledge/reasoning capabilities of $a$ over outputs that respect the constraint. This property is not true for alternative models, even if they only produce outputs respecting the constraint, and is a clear advantage of $g$ as a target model.
B.2 On the limitations of autoregressive models in representing $g$
@@
First, we come back to the situation described in the main text where the constraint $b$ prohibits the token “bad” to appear in $y$. We saw that fine-tuning could not completely guarantee this constraint, because it would never put zero mass on the token “bad”. Furthermore, fine-tuning would be pretty inefficient, even as an approximator, for this “avoidance” situation: a huge dataset would probably be necessary for significantly reducing the probability of “bad”, because the fine-tuning process would only see examples not containing “bad”, with no explicit negative examples. As mentioned in the text, another approach, not involving fine-tuning, would be, at each timestep $t$ during the generation of $y$, to remove the token “bad” from the softmax vector $p(\cdot|y_{<t})$, and renormalize this vector to sum to 1.111111This is a special case of a more general technique based on lexical constraints described by finite-state automata, see e.g. Willard and Louf [45]. This would result in an autoregressive model $a^{\prime}$ that always respects the constraint, but whose distribution would differ from $g$. Intuitively, this would be because the distribution for $g$, which corresponds to natural sequences from $a$ not containing “bad”, would anticipate early on, for the initial tokens of $y$, the fact that “bad” cannot appear later, while $a^{\prime}$ would ignore any such influence. [^12] The intuition may be clearer with a more situation-specific word such as “malaria”: in order to eliminate sequences containing this word from $a$, $g$ has to depreciate the probability of prefixes that naturally lead to that word — such as “This proliferation of mosquitoes raises the risk of” — while an $a^{\prime}$ intervening on the local generation of “malaria” would not perform such anticipation.
@@
Conversely, rather than trying to avoid a certain token such as “bad”, we could try to enforce the presence of a token such as “good” somewhere in the sequence $y$. In other words, we would have $b(y)=1$ iff $y$ contains the token “good”. Here, we could again sample a large dataset from $a$, keep only the sequences containing “good”, and fine-tune an ARM $a^{\prime}$ on the resulting dataset. However, while $a^{\prime}$ would be better than $a$ at producing sequences containing “good”, it could still produce other sequences, and therefore would not strictly guarantee the constraint (although this fine-tuning would probably be more effective as an approximator than in the previous example of avoiding “bad”, because at least it could rely on instances explicitly containing the token). Again, we could try another autoregressive approach: we could, at each step of the generation of $y$, check whether we have already produced the token “good”, and if not, continue the generation of $y$ (i.e., prevent the generation of an $\langle eos\rangle$ token). This would result in an ARM $a^{\prime}$, which would guarantee strict satisfaction of the constraint, but the distribution of the sequences produced by $a^{\prime}$ would be very different from $g$. This is because, similarly to the previous case, while $g$ would anticipate in its distribution over early tokens the necessity to produce “good” at a later stage, $a^{\prime}$ would totally ignore such influence of “good” over the early tokens. We intuitively expect that $a^{\prime}$ would tend to produce longer sequences than $g$, because there would be little incentive to produce “good” early on during the generation. [^13] Again, the intuition might be clearer if the constraint was about producing a rarer word than “good”, associated with a specific situation, for instance “trophy”. Sequences from $a$ containing this word, which would be selected by $g$, would tend to start differently than generic sequences from $a$, but $a^{\prime}$ would not show this kind of influence. This was verified through a simple experiment in App. E (Comparison to a heuristic sampler), which confirmed that such an approach does not provide a satisfactory approximation to $g$.
@@
For these two examples, the situation with standard LLMs such as $a$, built over neural networks, contrasts strongly with what would be the case if $a$ were an $n$-gram language model. In that case, $a$ would be equivalent to a probabilistic weighted finite-state automaton (WFSA) [32], a classical form of autoregressive model, and such constraints as the presence or absence of a specific token could be realized through intersection between such automata, resulting in another probabilistic automaton, with the same distribution as $g$. However, such intersections, which rely on dynamic programming over the finite set of states of these models, are unavailable in the domain of neural LMs. [^14] Zhang et al. [52] recently proposed a technique with some similarities to what we are describing here, by performing the intersection of an HMM (a formalism with strong connections with WFSAs) with a finite-state automaton imposing lexical constraints. In order to apply this technique to a standard neural LLM, the authors propose to first approximate this LLM with an HMM, before applying the finite-state constraints. See Section 5 and App. A for more details on that work.
Thus, to summarize, if $a$ were an $n$-gram language model, where $b$ was a constraint similar to those considered in the two examples above, the problem of modeling $g$ through an ARM would be solvable. Another instance where this problem is solvable — staying this time in the familiar domain of neural language models — is when $a$ is a standard LLM, and $b$ is the constraint which imposes the presence of a certain prefix $[y_{1};\ldots;y_{t}]$ at the beginning of the generated sequence $y$. It is easy to see that the generative process that samples $y_{t+1}\sim a(\cdot|[y_{1};\ldots;y_{t}])$, $y_{t+2}\sim a(\cdot|[y_{1};\ldots;y_{t+1}])$, and so on, in the standard way, corresponds to an autoregressive model $a^{\prime}$ that has the same distribution as $g$. However, again, this is a very special case, and as soon as the constraint does not refer to such an initial section of $y$, the problem becomes much more challenging.
@@
B.3 Details about Theorem 1
@@
Under the assumption $P\neq NP$, there exists a polynomial-time computable (PTC) ARM $a$ and a PTC binary predicate $b$ such that no PTC ARM $a^{\prime}$ has the same support [^15] The support of a distribution $p$ over ${\mathcal{Y}}$ is the set of $y$’s such that $p(y)>0$. as the gold model $g$, and in particular, has the same distribution as $g$.
@@ B.3.1 Proof sketch
Two examples of such sequence encodings would be: [^16] Here the token vocabulary would be $V=\{x,0,1,2,3,4,5,6,7,8,9,\vee,\neg,{;},\#\}$.
@@
B.4 Characterization of $g^{\prime}$
Let $P^{\prime}(y)\doteq a^{\prime}(y)b(y)$ and $p^{\prime}(y)$ be the distribution such that $p^{\prime}(y)\propto P^{\prime}(y)$. Then $p^{\prime}(y)=\frac{1}{Z^{\prime}}P^{\prime}(y)$ with $Z^{\prime}\doteq\sum_{y\in{\mathcal{Y}}}P^{\prime}(y)$. We have $P^{\prime}(y)\leq a^{\prime}(y)\cdot 1$, and $r(y)\doteq\frac{P^{\prime}(y)}{a^{\prime}(y)\cdot 1}=b(y)$. According to the fundamental result about rejection sampling, see [25, Section 2.2], $p^{\prime}$ is the distribution obtained by sampling $y\sim a^{\prime}$ and accepting $y$ with probability $r(y)$, which in our case is either 00 or $1$. This is exactly what Algorithm 3 does, and therefore $g^{\prime}$ has the same distribution as $p^{\prime}$, in other words: